# Sampling by averaging: A multiscale approach to score estimation

### v1

@@ 1 Introduction
Efficiently sampling from complex unnormalised probability distributions is a fundamental challenge across many scientific domains, including statistics, chemistry, computational physics, and biology, see, e.g. Lelièvre et al. [2010]; Liu [2008]; Gelman et al. [2013]. Classical Markov Chain Monte Carlo (MCMC) methods provide asymptotically unbiased samples under mild assumptions on the target [Durmus and Moulines, 2017; Mousavi-Hosseini et al., 2023]. However, they often become computationally inefficient in practice, especially for high-dimensional, multimodal targets, due to the need for long Markov chains to ensure adequate mixing and convergence. To alleviate these issues, annealing-based strategies construct a sequence of smoother intermediate distributions bridging a simple base distribution and the target. This principle underlies algorithms such as Parallel Tempering [Geyer and Thompson, 1995; Swendsen and Wang, 1986], Annealed Importance Sampling [Neal, 2001], and Sequential Monte Carlo [Doucet et al., 2001; Del Moral et al., 2006].
Recently, score-based diffusion methods [Hyvärinen, 2005; Song and Ermon, 2020] have emerged as powerful models capable of generating high-quality samples. These approaches simulate a reverse diffusion process guided by time-dependent score functions. While successful in generative modelling, they rely on access to training samples—a setting that differs from sampling problems, where only the unnormalised target density is available. Therefore, traditional score-matching techniques are not applicable. To address this, a growing body of work has explored ways to estimate the score function using only the target density, either through neural networks using variational objectives [Vargas et al., 2024; Richter and Berner, 2024; Chen et al., 2025] or training-free alternatives based on MCMC [Grenioux et al., 2024; Huang et al., 2024; Chen et al., 2024; Saremi et al., 2024; Phillips et al., 2024]. In this work, we present a novel training-free sampling framework that eliminates the need for the inner MCMC loop required in previous works to approximate the score function. Our main contributions are as follows:
@@ 2.1 Diffusion paths
The reverse process in diffusion models consists in sampling along a path of probability distributions $(\mu_{t})_{t\in[0,1]}$, which starts at a simple distribution and ends at the target distribution $\pi\propto e^{-V_{\pi}}$ on $\mathbb{R}^{d}$. Following Chehab and Korba [2024], the intermediate distributions can be expressed as follows
@@
where * denotes the convolution operation, $\nu$ describes the base or noising distribution, and $\lambda_{t}$ is an increasing function called schedule, such that, $\lambda_{t}\in[0,1]$ and $\lambda_{1}=1$. We refer to the path $(\mu_{t})_{t\in[0,1]}$ in (1) as the diffusion path. By choosing $\lambda_{t}=e^{-2T(1-t)}$ for some fixed $T$, we recover the path corresponding to a forward OU noising process, a widely used approach in diffusion models [Song et al., 2021; Benton et al., 2024; Chen et al., 2023]. In this case, the reverse-time dynamics can be characterised [Anderson, 1982], requiring only to estimate the score functions $\nabla\log\mu_{t}$ along the path. However, for general diffusion paths, the reverse process cannot be described by a closed-form SDE due to intractable control terms in the drift, which can be estimated using neural networks as in Albergo et al. [2023]. Alternatively, sampling from the diffusion path can be achieved through annealed Langevin dynamics [Song and Ermon, 2019; Cordero-Encinar et al., 2025], which also relies on score estimation but avoids dealing with intractable drift terms.
The diffusion path has demonstrated very good performance in the generative modelling literature where the score of the intermediate distributions can be estimated empirically using score matching techniques [Song and Ermon, 2019]. In the context of sampling, however, estimating the score is more challenging, as samples from the data distribution are not available and instead we only have access to an unnormalised target probability distribution. In particular, the expression of the score of the intermediate distributions is given by
@@
where $\rho_{t,x}(y)\propto\nu\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)e^{-V_{\pi}(y/\sqrt{\lambda_{t}})}$. Notably, in the limits $\lambda_{t}=0$ and $\lambda_{t}=1$, the conditional distribution $\rho_{t,x}$ converges to a Dirac delta function. The score in (2) involves an expectation over $\rho_{t,x}$, which typically requires MCMC sampling and results in a computationally expensive nested-loop structure. We avoid this by exploiting multiscale dynamics, as described in the following section.
@@ 2.2 Multiscale dynamics and stochastic averaging
$\mathrm{d}\bar{X}_{t}=\bar{b}(t,\bar{X}_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}t,\quad\quad\bar{b}(t,x)=\mathbb{E}_{Y\sim\rho_{t,x}}[b(t,x,Y)],$
where $\rho_{t,x}$ is the invariant measure of the frozen process $\mathrm{d}Y_{s}=f(t,x,Y_{s})\mathrm{d}s+\sqrt{2}\mathrm{d}\tilde{B}_{s}$. In computational statistics, stochastic slow-fast systems have been leveraged to simulate the averaged dynamics efficiently [Akyildiz et al., 2024; Weinan et al., 2005; Pavliotis and Stuart, 2008; Harvey et al., 2011]. In our context of sampling from diffusion paths, the averaged drift $\bar{b}(t,x)$ corresponds to the score $\nabla\log\mu_{t}(x)$, which is defined as an expectation (see Eq. (2)). By leveraging multiscale dynamics, we can approximate this expectation without relying on a nested-loop sampling structure—a key insight that underpins our method, as summarised in Section 3.
@@ 2.3 Related work
Several recent works have proposed sampling algorithms based on the diffusion path. We divide them into two categories: neural samplers and learning-free samplers. Neural samplers such as those in Vargas et al. [2024]; Richter and Berner [2024]; Chen et al. [2025]; Zhang and Chen [2022]; Noble et al. [2025] estimate the time-dependent drift function of the diffusion process using a parameterised model—typically a neural network—by solving a variational inference problem defined over the space of path measures. As in diffusion generative models, their performance is limited by the expressiveness of the chosen model class. In contrast, learning-free samplers [Huang et al., 2024; Chen et al., 2024; Grenioux et al., 2024; Saremi et al., 2024; He et al., 2024b; Phillips et al., 2024] do not require any training similar to our work. Instead, they estimate the score function $\nabla\log\mu_{t}$ using MCMC samples from the conditional distribution $\rho_{t,x}$. That is, these approaches involve a bi-level structure, with an inner MCMC loop used at each step of the outer diffusion-based sampler. In this sense, they implement a Langevin-within-Langevin strategy. Our approach avoids nested loops by leveraging multiscale dynamics along the diffusion path. This results in a Langevin-by-Langevin sampler that is both simpler and more computationally efficient.
Besides, while acceleration techniques based on underdamped dynamics [Duncan et al., 2017; Monmarché, 2020] have shown promise in the generative modelling literature [Dockhorn et al., 2022a, b; Holderrieth et al., 2024], and recent work by Blessing et al. [2025] introduces neural underdamped diffusion samplers, no training-free accelerated samplers based on diffusion processes have been proposed. Addressing this gap is one of the key contributions of our work.
@@
Our approach is also closely related to the class of proximal samplers [Chen et al., 2022; Lee et al., 2021b]. These methods define an auxiliary joint distribution of the form $\tilde{\pi}(x,z)\propto\pi(x)\exp(-\|x-z\|^{2}/(2\eta))$ and perform Gibbs sampling by alternating between the convolutional distribution $\tilde{\pi}(z|x)$ and the denoising posterior $\tilde{\pi}(x|z)$. In practice, sampling from the denoising posterior $\tilde{\pi}(x|z)$ involves finding a mode using gradient-based optimisation techniques, followed by rejection sampling around that mode. This step can be computationally expensive and inefficient, especially in high-dimensional settings. Our method improves upon these approaches by replacing the proximal-point-style posterior sampling step with a fast-timescale dynamics, enabling more efficient exploration of the denoising posterior. Furthermore, unlike proximal samplers that operate at a fixed noise level, our approach targets the entire diffusion path, smoothly interpolating between the base distribution and the target.
@@ 3 Sampling using multiscale dynamics
A key challenge in sampling from the diffusion path in (1) is accurately estimating the score function when we do not have direct access to samples. Since the score of the intermediate distributions along the diffusion path is computed as an expectation over the conditional distribution $\rho_{t,x}$ (see Eq. (2)), we propose to sample from the time-dependent conditional distribution $\rho_{t,x}$ using an alternative diffusion process with a shorter timescale than the original dynamics following the path in (1). This enables both processes to run in parallel, unlike existing methods that execute them sequentially. In those approaches, each iteration of the original diffusion process (outer loop) requires multiple time steps of the process targeting the conditional distribution $\rho_{t,x}$ (inner loop) to learn the score function, leading to increased computational cost and a higher number of evaluations of the target potential. By employing different timescales for each diffusion process, our method improves sampling efficiency while ensuring convergence to the correct target distribution, as justified by stochastic averaging theory [Liu et al., 2020]. We explore this approach in two settings: (i) when the original diffusion is a Langevin dynamics driven by the scores of the intermediate probability distribution $\nabla\log\mu_{t}$ (Section 3.1), and (ii) when we use the reverse SDE corresponding to an OU noising process (Section 3.2).
@@ 3.1 MultALMC: Multiscale Annealed Langevin Monte Carlo
Following Song and Ermon [2019]; Cordero-Encinar et al. [2025], we use a time inhomogeneous Langevin SDE to sample from the diffusion path (1) given by
@@
This system converges to the averaged dynamics described in (3) as $\varepsilon\rightarrow 0$, since taking the expectation of the drift term in the $X_{t}$ dynamics with respect to $Y_{t}\sim\hat{\rho}_{t,X_{t}}$ recovers the score function $\nabla\log\hat{\mu}_{t}$ (see Equation (2) and Section 4). For numerical implementation, we use an adaptive scheme that alternates between the two expressions for $X_{t}$ depending on the value of $\lambda_{\kappa t}$ in order to avoid numerical instabilities as $\lambda_{\kappa t}$ approaches 0 or $1$. The resulting system of SDEs can be discretised using different numerical schemes, however, standard integrators such as Euler-Maruyama (EM) will be unstable when $\varepsilon$ is small, due to the large-scale separation between the slow and fast processes. To mitigate this, we leverage discretisations which remain stable, despite the stiffness of the SDE, such as the exponential integrator [Hochbruck and Ostermann, 2010] and the SROCK method [Abdulle et al., 2018]. Furthermore, since $\rho_{t,x}$ converges to a Dirac delta distribution centred at 0 when $\lambda_{\kappa t}=0$ and centred at $x$ when $\lambda_{\kappa t}=1$, we apply a further slight modification to the slow-fast system to improve numerical stability (see App. B.1.1 for details).
@@
When implemented in practice, the overdamped system in (4) requires a very small value of $\kappa$ (which corresponds to a slowly varying dynamics driven by $\nabla\log\hat{\mu}_{t}$) to accurately follow the marginals of the path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$ and ultimately sample from the target distribution. This results in a large number of discretisation steps, thus making the method computationally expensive, see App. E.1 for more details. In contrast, underdamped dynamics exhibit faster mixing times [Bou-Rabee and Eberle, 2023; Eberle and Lörler, 2024], which intuitively helps the system explore the space more efficiently and better track the intermediate distributions along the path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$. These dynamics have also shown strong empirical performance in the neural sampling literature [Blessing et al., 2025], requiring fewer discretisation steps. Motivated by these advantages, we propose augmenting the state space with auxiliary velocity variables $\bar{V}_{t}$, and adopting the following underdamped Langevin diffusion to sample from the path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$
@@
The mass parameter $M$ determines the coupling between position ${X}_{t}$ and velocity ${V}_{t}$, while the friction coefficient $\Gamma$ controls the strength of noise injected into the velocity component. Both $M$ and $\Gamma$ are symmetric positive definite matrices. The behaviour of the system is governed by the interplay between $M$ and $\Gamma$ [McCall, 2010]. Based on this, we propose to use the following underdamped slow-fast system to sample from the target distribution
@@
Note that in the underdamped setting, both $X_{t},V_{t}$ are treated as slow variables, and as $\varepsilon\to 0$, this system converges to the averaged dynamics in (5). To implement the system in practice, we combine a symmetric splitting scheme (OBABO) [Monmarché, 2020] for the slow-dynamics, and discretise the fast-dynamics using SROCK method [Abdulle et al., 2018]. We will evaluate the performance of the proposed sampler based on this underdamped slow-fast system (Algorithm 2)—referred to as MultALMC—under two choices of the reference distribution $\nu$: a standard Gaussian and a Student’s t distribution. A more detailed description of the implementation is provided in App. B.1.2.
@@
When the target distribution has heavy tails, standard Gaussian diffusions often fail to capture the correct tail behaviour [Pandey et al., 2025; Shariatian et al., 2025]. This motivates the use of heavy-tailed noising processes, which have been shown to offer theoretical guarantees in such settings [Cordero-Encinar et al., 2025].
We propose sampling along a heavy-tailed diffusion path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$, where the reference distribution $\nu$ is chosen as a Student’s t-distribution with tail index $\alpha$, i.e., $\nu\propto(1+\|x\|^{2}/\alpha)^{-(\alpha+d)/2}$. This can be implemented using the underdamped slow-fast system defined in (6). However, when the target distribution is heavy-tailed, the use of overdamped Langevin dynamics for the fast component results in slow convergence [Mousavi-Hosseini et al., 2023; Wang, 2006]. This motivates the consideration of alternative diffusion processes for the fast dynamics that mix more efficiently under such conditions. Specifically, we explore a modified Itô diffusion proposed in He et al. [2024a], defined as
@@
This modified Itô diffusion for heavy-tailed distributions is shown to have faster convergence guarantees when the target has finite variance [He et al., 2024a]. By replacing the fast process in (6) with the diffusion defined in (7), we obtain samplers better suited for heavy-tailed targets, with improved convergence properties.

### v2

@@ 1 Introduction
Efficiently sampling from complex unnormalised probability distributions is a fundamental challenge across many scientific domains, including statistics, chemistry, computational physics, and biology, see, e.g. Lelièvre et al. (2010); Liu (2008); Gelman et al. (2013). Classical Markov Chain Monte Carlo (MCMC) methods provide asymptotically unbiased samples under mild assumptions on the target (Durmus and Moulines, 2017; Mousavi-Hosseini et al., 2023). However, they often become computationally inefficient in practice, especially for high-dimensional, multimodal targets, due to the need for long Markov chains to ensure adequate mixing and convergence. To alleviate these issues, annealing-based strategies construct a sequence of smoother intermediate distributions bridging a simple base distribution and the target. This principle underlies algorithms such as Parallel Tempering (Geyer and Thompson, 1995; Swendsen and Wang, 1986), Annealed Importance Sampling (Neal, 2001), and Sequential Monte Carlo (Doucet et al., 2001; Del Moral et al., 2006).
Recently, score-based diffusion methods (Hyvärinen, 2005; Song and Ermon, 2020) have emerged as powerful models capable of generating high-quality samples. These approaches simulate a reverse diffusion process guided by time-dependent score functions. While successful in generative modelling, they rely on access to training samples—a setting that differs from sampling problems, where only the unnormalised target density is available. Therefore, traditional score-matching techniques are not applicable. To address this, a growing body of work has explored ways to estimate the score function using only the target density, either through neural networks using variational objectives (Vargas et al., 2024; Richter and Berner, 2024; Chen et al., 2025) or training-free alternatives based on MCMC (Grenioux et al., 2024; Huang et al., 2024; Chen et al., 2024; Saremi et al., 2024; Phillips et al., 2024). In this work, we present a novel training-free sampling framework that eliminates the need for the inner MCMC loop required in previous works to approximate the score function. Our main contributions are as follows:
@@ 2.1 Diffusion paths
The reverse process in diffusion models consists in sampling along a path of probability distributions $(\mu_{t})_{t\in[0,1]}$, which starts at a simple distribution and ends at the target distribution $\pi\propto e^{-V_{\pi}}$ on $\mathbb{R}^{d}$. Following Chehab and Korba (2024), the intermediate distributions can be expressed as follows
@@
where * denotes the convolution operation, $\nu$ describes the base or noising distribution, and $\lambda_{t}$ is an increasing function called schedule, such that, $\lambda_{t}\in[0,1]$ and $\lambda_{1}=1$. We refer to the path $(\mu_{t})_{t\in[0,1]}$ in (1) as the diffusion path. By choosing $\lambda_{t}=e^{-2T(1-t)}$ for some fixed $T$, we recover the path corresponding to a forward OU noising process, a widely used approach in diffusion models (Song et al., 2021; Benton et al., 2024; Chen et al., 2023). In this case, the reverse-time dynamics can be characterised (Anderson, 1982), requiring only to estimate the score functions $\nabla\log\mu_{t}$ along the path. However, for general diffusion paths, the reverse process cannot be described by a closed-form SDE due to intractable control terms in the drift, which can be estimated using neural networks as in Albergo et al. (2023). Alternatively, sampling from the diffusion path can be achieved through annealed Langevin dynamics (Song and Ermon, 2019; Cordero-Encinar et al., 2025), which also relies on score estimation but avoids dealing with intractable drift terms.
The diffusion path has demonstrated very good performance in the generative modelling literature where the score of the intermediate distributions can be estimated empirically using score matching techniques (Song and Ermon, 2019). In the context of sampling, however, estimating the score is more challenging, as samples from the data distribution are not available and instead we only have access to an unnormalised target probability distribution. In particular, under the assumption that $\nu,\pi$ are bounded and have finite second-order moments, and that $|\nabla\log\nu|$ and $|\nabla V_{\pi}|$ are bounded, the expression of the score of the intermediate distributions is given by
@@
where $\rho_{t,x}(y)\propto\nu\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)e^{-V_{\pi}(y/\sqrt{\lambda_{t}})}$ and we have used that $\nabla(f*g)=(\nabla f)*g=f*(\nabla g)$ when $f$ and $g$ are differentiable functions. Notably, in the limits $\lambda_{t}=0$ and $\lambda_{t}=1$, the conditional distribution $\rho_{t,x}$ converges to a Dirac delta function. The score in (2) involves an expectation over $\rho_{t,x}$, which typically requires MCMC sampling and results in a computationally expensive nested-loop structure. We avoid this by exploiting multiscale dynamics, as described in the following section.
@@ 2.2 Multiscale dynamics and stochastic averaging
$\mathrm{d}\bar{X}_{t}=\bar{b}(t,\bar{X}_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},\quad\quad\bar{b}(t,x)=\mathbb{E}_{Y\sim\rho_{t,x}}[b(t,x,Y)],$
where $\rho_{t,x}$ is the invariant measure of the frozen process $\mathrm{d}Y_{s}=f(t,x,Y_{s})\mathrm{d}s+\sqrt{2}\mathrm{d}\tilde{B}_{s}$. In computational statistics, stochastic slow-fast systems have been leveraged to simulate the averaged dynamics efficiently (Akyildiz et al., 2024; Weinan et al., 2005; Pavliotis and Stuart, 2008; Harvey et al., 2011). In our context of sampling from diffusion paths, the averaged drift $\bar{b}(t,x)$ corresponds to the score $\nabla\log\mu_{t}(x)$, which is defined as an expectation (see Eq. (2)). By leveraging multiscale dynamics, we can approximate this expectation without relying on a nested-loop sampling structure, a key insight that underpins our method, as summarised in Section 3.
@@ 2.3 Related work
Several recent works have proposed sampling algorithms based on the diffusion path. We divide them into two categories: neural samplers and learning-free samplers. Neural samplers such as those in Vargas et al. (2024); Richter and Berner (2024); Chen et al. (2025); Zhang and Chen (2022); Noble et al. (2025) estimate the time-dependent drift function of the diffusion process using a parametrised model, typically a neural network, by solving a variational inference problem defined over the space of path measures. As in diffusion generative models, their performance is limited by the expressiveness of the chosen model class. In contrast, learning-free samplers (Huang et al., 2024; Chen et al., 2024; Grenioux et al., 2024; Saremi et al., 2024; He et al., 2024b; Phillips et al., 2024) do not require any training similar to our work. Instead, they estimate the score function $\nabla\log\mu_{t}$ using MCMC samples from the conditional distribution $\rho_{t,x}$. That is, these approaches involve a bi-level structure, with an inner MCMC loop used at each step of the outer diffusion-based sampler. In this sense, they implement a Langevin-within-Langevin strategy. Our approach avoids nested loops by leveraging multiscale dynamics along the diffusion path. This results in a Langevin-by-Langevin sampler that is both simpler and more computationally efficient.
Besides, while acceleration techniques based on underdamped dynamics (Duncan et al., 2017; Monmarché, 2020) have shown promise in the generative modelling literature (Dockhorn et al., 2022a, b; Holderrieth et al., 2024), and recent work by Blessing et al. (2025) introduces neural underdamped diffusion samplers, no training-free accelerated samplers based on diffusion processes have been proposed. Addressing this gap is one of the key contributions of our work.
@@
Our approach is also closely related to the class of proximal samplers (Chen et al., 2022; Lee et al., 2021b). These methods define an auxiliary joint distribution of the form $\tilde{\pi}(x,z)\propto\pi(x)\exp(-\|x-z\|^{2}/(2\eta))$ and perform Gibbs sampling by alternating between the convolutional distribution $\tilde{\pi}(z|x)$ and the denoising posterior $\tilde{\pi}(x|z)$. In practice, sampling from the denoising posterior $\tilde{\pi}(x|z)$ involves finding a mode using gradient-based optimisation techniques, followed by rejection sampling around that mode. This step can be computationally expensive and inefficient, especially in high-dimensional settings. Our method improves upon these approaches by replacing the proximal-point-style posterior sampling step with a fast-timescale dynamics, enabling more efficient exploration of the denoising posterior. Furthermore, unlike proximal samplers that operate at a fixed noise level, our approach targets the entire diffusion path, smoothly interpolating between the base distribution and the target.
@@ 3 Sampling using multiscale dynamics
A key challenge in sampling from the diffusion path in (1) is accurately estimating the score function when we do not have direct access to samples. Since the score of the intermediate distributions along the diffusion path is computed as an expectation over the conditional distribution $\rho_{t,x}$ (see Eq. (2)), we propose to sample from the time-dependent conditional distribution $\rho_{t,x}$ using an alternative diffusion process with a shorter timescale than the original dynamics following the path in (1). This enables both processes to run in parallel, unlike existing methods that execute them sequentially. In those approaches, each iteration of the original diffusion process (outer loop) requires multiple time steps of the process targeting the conditional distribution $\rho_{t,x}$ (inner loop) to learn the score function, leading to increased computational cost and a higher number of evaluations of the target potential. By employing different timescales for each diffusion process, our method improves sampling efficiency while ensuring convergence to the correct target distribution, as justified by stochastic averaging theory (Liu et al., 2020). We explore this approach in two settings: (i) when the original diffusion is a Langevin dynamics driven by the scores of the intermediate probability distribution $\nabla\log\mu_{t}$ (Section 3.1), and (ii) when we use the reverse SDE corresponding to an OU noising process (Section 3.2).
@@ 3.1 MultALMC: Multiscale Annealed Langevin Monte Carlo
Following Song and Ermon (2019); Cordero-Encinar et al. (2025), we use a time inhomogeneous Langevin SDE to sample from the diffusion path (1) given by
@@
This system converges to the averaged dynamics described in (3) as $\varepsilon\rightarrow 0$, since taking the expectation of the drift term in the $X_{t}$ dynamics with respect to $Y_{t}\sim\hat{\rho}_{t,X_{t}}$ recovers the score function $\nabla\log\hat{\mu}_{t}$ (see Equation (2) and Section 4). For numerical implementation, we use an adaptive scheme that alternates between the two expressions for $X_{t}$ depending on the value of $\lambda_{\kappa t}$ in order to avoid numerical instabilities as $\lambda_{\kappa t}$ approaches 0 or $1$. The resulting system of SDEs can be discretised using different numerical schemes, however, standard integrators such as Euler-Maruyama (EM) will be unstable when $\varepsilon$ is small, due to the large-scale separation between the slow and fast processes. To mitigate this, we leverage discretisations which remain stable, despite the stiffness of the SDE, such as the exponential integrator (Hochbruck and Ostermann, 2010) and the SROCK method (Abdulle et al., 2018). Furthermore, since $\rho_{t,x}$ converges to a Dirac delta distribution centred at 0 when $\lambda_{\kappa t}=0$ and centred at $x$ when $\lambda_{\kappa t}=1$, we apply a further slight modification to the slow-fast system to improve numerical stability (see App. B.1.1 for details).
@@
When implemented in practice, the overdamped system in (4) requires a very small value of $\kappa$ (which corresponds to a slowly varying dynamics driven by $\nabla\log\hat{\mu}_{t}$) to accurately follow the marginals of the path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$ and ultimately sample from the target distribution. This results in a large number of discretisation steps, thus making the method computationally expensive, see App. E.1 for more details. In contrast, underdamped dynamics exhibit faster mixing times (Bou-Rabee and Eberle, 2023; Eberle and Lörler, 2024), which intuitively helps the system explore the space more efficiently and better track the intermediate distributions along the path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$. These dynamics have also shown strong empirical performance in the neural sampling literature (Blessing et al., 2025), requiring fewer discretisation steps. Motivated by these advantages, we propose augmenting the state space with auxiliary velocity variables $\bar{V}_{t}$, and adopting the following underdamped Langevin diffusion to sample from the path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$
@@
The mass parameter $M$ determines the coupling between position ${X}_{t}$ and velocity ${V}_{t}$, while the friction coefficient $\Gamma$ controls the strength of noise injected into the velocity component. Both $M$ and $\Gamma$ are symmetric positive definite matrices, in practice we consider $M=mId$ and $\Gamma=\gamma Id$. The behaviour of the system is governed by the interplay between $M$ and $\Gamma$ (McCall, 2010). Based on this, we propose to use the following underdamped slow-fast system to sample from the target distribution
@@
Note that in the underdamped setting, both $X_{t},V_{t}$ are treated as slow variables, and as $\varepsilon\to 0$, this system converges to the averaged dynamics in (5). To implement the system in practice, we combine a symmetric splitting scheme (OBABO) (Monmarché, 2020) for the slow-dynamics, and discretise the fast-dynamics using SROCK method (Abdulle et al., 2018). We will evaluate the performance of the proposed sampler based on this underdamped slow-fast system (Algorithm 3), referred to as MultALMC, under two choices of the reference distribution $\nu$: a standard Gaussian and a Student’s t distribution. A simplified version of the implementation is presented in Alg. 1, while a more detailed description is provided in App. B.1.2.
Algorithm 1 MultALMC sampler: accelerated version
Schedule function $\lambda_{t}$, value for $\lambda_{\delta}$ and $\tilde{\lambda}$, friction coefficient $\Gamma$, mass parameter $M$, number of sampling steps $L$, time discretisation $0=T_{0}<\dots<T_{L}=1/\kappa$, step size $h_{l}=T_{l+1}-T_{l}$. Constants $\bm{\mu},\bm{\nu},\bm{\kappa}$ for SROCK step from (14), (15).
Initial samples ${X}_{0}\sim\mathcal{N}(0,I),{V}_{0}\sim\mathcal{N}(0,MI),Y_{0}\sim\mathcal{N}(0,I)$, $\lambda_{\kappa t}^{\prime}=\begin{cases}\lambda_{\delta}&\text{if $0\leq\lambda_{\kappa t}<\lambda_{\delta}$}\\ \lambda_{\kappa t}&\text{if $\lambda_{\delta}\leq\lambda_{\kappa t}\leq 1$}.\end{cases}$
for $l=0$ to $L$ do
  $\xi_{l}^{(1)},\xi_{l}^{(2)},\xi_{l}^{(3)}\sim\mathcal{N}(0,I)$
  if $0\leq\lambda_{\kappa T_{l}}<1-\lambda_{\delta}$ then
Half-step for velocity component
$V_{l}^{\prime}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}+\sqrt{h_{l}\Gamma}\xi_{l}^{(1)}\,,\qquad\qquad V_{l}^{\prime\prime}=\begin{cases}V_{l}^{\prime}+\frac{h_{l}}{2\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l}-Y_{l}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\\ V_{l}^{\prime}-\frac{h_{l}}{2\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\end{cases}$
$\qquad\qquad\qquad\texttt{Full EM-step for position component}\qquad X_{l+1}=X_{l}+h_{l}M^{-1}V_{l}^{\prime\prime}$
$\qquad\qquad\qquad\texttt{Full SROCK-step for fast component}\qquad Y_{l+1}=f(Y_{l},X_{l+1},\lambda_{\kappa T_{l}}^{\prime},h_{l},\varepsilon,\bm{\mu},\bm{\nu},\bm{\kappa})+\sqrt{\frac{2h}{\varepsilon}}\xi_{l}^{(2)}$
Half-step for velocity component
$V_{l}^{\prime\prime\prime}=\begin{cases}V_{l}^{\prime\prime}+\frac{h_{l}}{2\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l+1}-Y_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\,,\\ V_{l}^{\prime\prime}-\frac{h_{l}}{2\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l+1}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\,,\end{cases},\qquad V_{l+1}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}^{\prime\prime\prime}+\sqrt{h_{l}\Gamma}\xi_{l}^{(3)}$
  if $1-\lambda_{\delta}\leq\lambda_{\kappa T_{l}}\leq 1$ then
   Half-step for velocity component
   $\quad V_{l}^{\prime}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}+\sqrt{h_{l}\Gamma}\xi_{l}^{(1)}\,,\quad\quad V_{l}^{\prime\prime}=V_{l}^{\prime}-\frac{h_{l}}{2}\nabla V_{\pi}\left({X_{l}}\right)$
   Full EM-step for position component $\quad X_{l+1}=X_{l}+h_{l}M^{-1}V_{l}^{\prime\prime}$
   Half-step for velocity component
   $\quad V_{l}^{\prime\prime\prime}=V_{l}^{\prime\prime}-\frac{h_{l}}{2}\nabla V_{\pi}\left(X_{l+1}\right)\,,\quad\quad V_{l+1}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}^{\prime\prime\prime}+\sqrt{h_{l}\Gamma}\xi_{l}^{(3)}$   end for return $(X_{L},V_{L})$
@@
When the target distribution has heavy tails, standard Gaussian diffusions often fail to capture the correct tail behaviour (Pandey et al., 2025; Shariatian et al., 2025). This motivates the use of heavy-tailed noising processes, which have been shown to offer theoretical guarantees in such settings (Cordero-Encinar et al., 2025).
We propose sampling along a heavy-tailed diffusion path $(\hat{\mu}_{t})_{t\in[0,1/\kappa]}$, where the reference distribution $\nu$ is chosen as a Student’s t-distribution with tail index $\alpha$, i.e., $\nu\propto(1+\|x\|^{2}/\alpha)^{-(\alpha+d)/2}$. This can be implemented using the underdamped slow-fast system defined in (6). However, when the target distribution is heavy-tailed, the use of overdamped Langevin dynamics for the fast component results in slow convergence (Mousavi-Hosseini et al., 2023; Wang, 2006). This motivates the consideration of alternative diffusion processes for the fast dynamics that mix more efficiently under such conditions. Specifically, we explore a modified Itô diffusion proposed in He et al. (2024a), defined as
@@
This modified Itô diffusion for heavy-tailed distributions is shown to have faster convergence guarantees when the target has finite variance (He et al., 2024a). By replacing the fast process in (6) with the diffusion defined in (7), we obtain samplers better suited for heavy-tailed targets, with improved convergence properties.